complete the ordered pair for the equation3x+4y=12 (0, ),( ,0),(-4, )
Answer by Maths68(1474)

3x+4y=12
(0, )
3(0)+4y=12
4y=12
y=12/4
y=3
(0,3)



( ,0)
3x+4y=12
3x+4(0)=12
3x=12
x=12/3
x=4
(4,0)


(-4, )
3(-4)+4y=12
-12+4y=12
4y=12+12
4y=24
y=24/4
y=6
(-4,6)